OpenStudy (anonymous):

Jill rides her horse 5 miles due north and then 3 miles due east. What is the magnitude and direction of Jill's ride? Round to the nearest tenth. (1 point)5 mi; 45° northeast 4 mi; 31° east of north 5.8 mi; 31° east of north 8 mi; 60° east of north

OpenStudy (anonymous):

Listen, if all your homework questions are like this, you can ask one or two questions here, but there's no point getting somebody else to do them ALL for you. Try one yourself, or ask somebody to help you. The answers are of little use in themselves.

OpenStudy (anonymous):

5.8 mi; 31° east of north

OpenStudy (anonymous):

 distance is AB. use pythag method to calculate AB i.e. AB =\[\sqrt{5^{2}+3^{2}} = \sqrt{34} = 5.8309518948453004708741528775456\] rounded to 5.8 use trig function to get that angle. opposie/adjacent = tan(x) i.e. tan(x) = 3/5 = 0.6 x = arctan(3/5) = 30.9637565321 rounded to 31 Hence tha above answer. Let me know if that is correct.